# Generalizable Error Modeling for Search Relevance Data Annotation Tasks

### v1

@@ Abstract
Human data annotation is critical in shaping the quality of machine learning (ML) and artificial intelligence (AI) systems. One significant challenge in this context is posed by annotation errors, as their effects can degrade the performance of ML models. This paper presents a predictive error model trained to detect potential errors in search relevance annotation tasks for three industry-scale ML applications (music streaming, video streaming, and mobile apps) and assesses its potential to enhance the quality and efficiency of the data annotation process. Drawing on real-world data from an extensive search relevance annotation program, we illustrate that errors can be predicted with moderate model performance (AUC=0.65-0.75) and that model performance generalizes well across applications (i.e., a global, task-agnostic model performs on par with task-specific models). We present model explainability analyses to identify which types of features are the main drivers of predictive performance. Additionally, we demonstrate the usefulness of the model in the context of auditing, where prioritizing tasks with high predicted error probabilities considerably increases the amount of corrected annotation errors (e.g., 40% efficiency gains for the music streaming application). These results underscore that automated error detection models can yield considerable improvements in the efficiency and quality of data annotation processes. Thus, our findings reveal critical insights into effective error management in the data annotation process, thereby contributing to the broader field of human-in-the-loop ML.
@@ 1 Introduction
One of the key components of modern machine learning (ML) and artificial intelligence (AI) systems is the ever-increasing volume of human-annotated data used to train, evaluate, monitor, and update such models. The importance of this subject can hardly be overstated, as the quality of ML models is linked to the quality of the data they are trained and evaluated on [1, 2, 3, 4, 5]. With the data annotation market reaching an estimated \$1 billion in 2023 [6], it is important to note that the impact of this industry extends beyond this figure. The revenue that is tied to the ML products relying on human annotations is far greater than the data annotation market itself, underscoring the economic influence and strategic relevance of human annotations. However, a significant challenge lies in the management of annotation errors and their downstream consequences. The effects of these errors can cause setbacks in ML applications, making their management a critical task in maintaining the overall quality and effectiveness of ML models [1].
@@
In supervised ML, data annotation errors can impact model quality in at least two ways: First, mislabeling of training data is one of the main sources of prediction errors, as it prevents models from picking up on the true target classification function [3, 4]. Second, mislabeling of testing data distorts model evaluations, therefore affecting model selection and leading to the deployment of suboptimal models [5]. For example, correcting labeling errors in the widely used ImageNet [7] benchmark dataset would change the rank order of model performance in favor of lower capacity models [5].
@@
Several potential solutions have been proposed to tackle annotation errors. These include developing clear annotation guidelines paired with extensive annotator training [8], generating crowd-sourced annotations and monitoring inter-annotator agreement [9], implementing audits, quality checks, or relabeling by subject matter experts [2], as well as automated error detection techniques [10, 2]. Generally, these practices should be implemented in a way that allows for feedback loops and iterative refinement. For example, detected errors can be used as an opportunity to refine guidelines and provide more training to annotators. Similarly, annotators should be able to ask questions or raise concerns about the annotation process or specific items to help identify problems early [11, 12].
@@
The present paper explores automated error models and their potential to improve the efficiency of the data annotation process, specifically with respect to the quality and turnaround of annotations. We further analyze how automated error detection models can improve related processes such as auditing and relabeling. Compared to previous work addressing annotation errors, our approach introduces a unique emphasis on task-agnostic error prediction. Unlike many prior models, which focus on specific tasks and require separate training for each labeling task, our models are trained to predict errors directly from a combination of task features and annotator features, offering a higher degree of generalizability.
@@ 1.1 Background and Related Work
Automated error detection techniques aim to identify instances that may be labeled incorrectly or inconsistently within a given dataset [10]. Once these potential errors are detected, they can be forwarded for manual review by human annotators or processed via annotation error correction methods. Error detection has been implemented across diverse domains in natural language processing [13, 14, 10], and image classification [5, 2]. However, the logic of error modeling can be applied to a variety of supervised ML tasks.
@@
One key distinction between error detection techniques is whether the approach is variation-based, rule-based, or model-based [10]. Variation-based approaches generally rely on inputs labeled by multiple annotators. For example, in settings where only one label can be correct, disagreement across annotators can indicate annotation errors. Similarly, annotator disagreement can be captured by computing variation metrics like label entropy across annotators, such that high entropy would be indicative of annotation errors [15]. While variation-based error detection methods are easy to implement, one major disadvantage is that they are limited to crowd-sourced scenarios, where more than one annotator label is produced for each instance. Requiring multiple labels can introduce significant costs to annotation processes.
@@
Rule-based techniques, on the other hand, are useful when there are clear annotation guidelines that can be translated into formal rules. In such situations, annotations that are impossible given the set of rules would be flagged as incorrect. For example, errors can be flagged if an annotator picks a label that is not applicable to the given task. In language annotation tasks, n-grams that are lexically or morphologically impossible would fall into this category [16]. However, while rule-based approaches can prove to be highly effective, they are typically tailored to a very specific task, imposing limits on generalizability. Additionally, they only apply to a specific subset of errors, namely those that are the result of objective rule violations.
@@
Model-based error detection approaches typically utilize probabilistic classifiers designed to detect annotation errors. Past work has frequently relied on the models predicting an annotation label (typically a multi-class classification problem) and subsequently comparing the predicted label to a label provided by a human annotator. If there is disagreement between the human annotation and the model-generated annotation, the annotation can be marked as inconsistent and additionally directed to a human auditor [17, 18]. Alternatively, model confidence scores or predicted probabilities can provide an indication of task difficulty [19]. Implicitly, these approaches rely on the assumption that models perform worse on instances that are more difficult for human annotators as well. Consequently, instances that show low model performance - either by deviation from human annotations or by low model confidence scores - would be more likely to contain annotation errors.
@@
However, it is important to note that a model’s confidence score on the labeling task itself may not always be a reliable indicator of a human annotator’s potential error. For instance, low model confidence could be due to the model’s limitations in representing complex, nuanced, or context-specific tasks rather than indicating an annotation error. Conversely, there may be cases where a model is highly confident and yet wrong due to biases in training data or overly simplistic decision boundaries. Therefore, while using model confidence scores can certainly be a useful tool in identifying potential annotation errors, it should be treated as part of a broader toolkit and used in conjunction with other methods. For a more comprehensive approach to error detection, it might be beneficial also to consider other factors such as the context of the task, the difficulty level of the task, and the reliability and expertise of the human annotators themselves. Relatedly, this approach can be less effective in identifying systematic errors, where both the model and the human annotators consistently misinterpret certain types of instances. Such errors could go undetected as model predictions and human annotations would agree, despite being incorrect. Finally, predicting labels may suffer from low generalizability, as a separate model needs to be trained for each specific labeling task, which can be computationally expensive and may not be feasible for all tasks, especially those requiring complex models.
@@
As an alternative approach explored in this paper, a model can be used to directly predict whether a human label is correct or not (typically a binary classification problem) without explicitly predicting the underlying label itself. The advantage of this approach is that it is often easier to predict the presence of an error compared to predicting the correct label. Relatedly, directly predicting the correctness of a label can be computationally more efficient, as it does not necessarily need to consider all possible label classes, which can be vast in certain tasks. Overall this approach makes it easier to train, evaluate, and maintain such models even when training data are limited. Additionally, error detection models can be set up to be task agnostic: a model predicting errors instead of labels may generalize to a wider range of tasks because certain features are applicable to many different scenarios. For example, taking into account general measures of task complexity, past annotator performance, time on task, or comments provided by annotators can be informative irrespective of the underlying labeling task. Past research has rarely applied this approach [10] since most studies are based on archival data for which no annotator data (e.g., past performance) or task-completion data (e.g., time on task) are available.
@@ 1.2 Current Research
The current research presents a study on behavioral error detection models and their applications for search relevance annotation tasks. These tasks were drawn from an extensive data annotation program that generates training and evaluation data for search and recommendation systems. Here, we focus on search relevance annotations related to music streaming, video streaming, and mobile applications. A detailed description of the annotation task can be found in the Method section below.
Drawing on authentic data from the search relevance annotation program, we first use model-based error detection techniques to show that annotation errors can be detected with moderate classification performance. We demonstrate the predictability of errors for different search relevance rating tasks for all three ML applications: music streaming, video streaming, and mobile applications. Second, using Shapley values [20], we explore the contributions of a variety of annotator and task-related features in order to better understand which sources of information are driving the model’s predictions. Third, we explore the generalizability of our error detection models across rating tasks and ML applications. Fourth, we show that our error detection models can unlock considerable efficiency gains for auditing and relabeling by prioritizing tasks with high predicted error probabilities at the expense of tasks with low predicted error probabilities. Taken together, our findings show that error detection models can be highly effective in improving the efficiency of data annotation processes.
@@ 2.2 Search Relevance Annotation Tasks
p6cmp9.7cm Feature Name Description
@@
   input_language Categorical variable distinguishing between 21 different input languages
   input_query_type Intent classification for query

### v2

# Generalizable Error Modeling for Human Data Annotation: Evidence From an Industry-Scale Search Data Annotation Program

@@ Abstract
Machine learning (ML) and artificial intelligence (AI) systems rely heavily on human-annotated data for training and evaluation. A major challenge in this context is the occurrence of annotation errors, as their effects can degrade model performance. This paper presents a predictive error model trained to detect potential errors in search relevance annotation tasks for three industry-scale ML applications (music streaming, video streaming, and mobile apps). Drawing on real-world data from an extensive search relevance annotation program, we demonstrate that errors can be predicted with moderate model performance (AUC=0.65-0.75) and that model performance generalizes well across applications (i.e., a global, task-agnostic model performs on par with task-specific models). In contrast to past research, which has often focused on predicting annotation labels from task-specific features, our model is trained to predict errors directly from a combination of task features and behavioral features derived from the annotation process, in order to achieve a high degree of generalizability. We demonstrate the usefulness of the model in the context of auditing, where prioritizing tasks with high predicted error probabilities considerably increases the amount of corrected annotation errors (e.g., 40% efficiency gains for the music streaming application). These results highlight that behavioral error detection models can yield considerable improvements in the efficiency and quality of data annotation processes. Our findings reveal critical insights into effective error management in the data annotation process, thereby contributing to the broader field of human-in-the-loop ML.
@@ 1 Introduction
One of the key components of modern machine learning (ML) and artificial intelligence (AI) systems is the ever-increasing volume of human-annotated data used to train, evaluate, monitor, and update such models. The importance of this subject can hardly be overstated, as the quality of ML models is closely linked to the quality of the data they are trained and evaluated on [sambasivan_everyone_2021, bernhardt_active_2022, amershi_modeltracker_2015, meek_characterization_2016, northcutt_pervasive_2021]. With the data annotation market reaching an estimated \$15 billion in 2023 [verified_market_research_data_2023], it is important to note that the impact of this industry extends far beyond this figure, as the revenue tied to the ML and AI products relying on human data annotation is much greater than the data annotation market itself. Therefore, a significant challenge lies in the management of annotation errors and their downstream consequences [sambasivan_everyone_2021].
@@
In supervised ML, data annotation errors can impact model quality in at least two ways: First, mislabeling of training data is one of the main sources of prediction errors, as it prevents models from picking up on the true target function [amershi_modeltracker_2015, meek_characterization_2016]. Second, mislabeling of testing data distorts model evaluations, therefore affecting model selection and leading to the deployment of suboptimal models [northcutt_pervasive_2021]. For example, correcting labeling errors in the widely used ImageNet [deng_imagenet_2009] benchmark dataset would change the rank order of model performance in favor of lower capacity models [northcutt_pervasive_2021].
@@
Several potential solutions have been proposed to tackle annotation errors. These include developing clear annotation guidelines paired with extensive annotator training [rottger_two_2022], generating crowd-sourced annotations and monitoring inter-annotator agreement [aroyo_truth_2015], implementing audits, quality checks, or relabeling by subject matter experts [bernhardt_active_2022], as well as automated error detection techniques [klie_annotation_2022, bernhardt_active_2022]. Generally, these practices should be implemented in a way that allows for feedback loops and iterative refinement. For example, detected errors can be used as an opportunity to refine guidelines and provide additional training to annotators. Similarly, annotators should be able to ask questions or raise concerns about the annotation process or specific items to help identify problems early [tseng_best_2020, mcdonnell_why_2016].
@@
The present paper explores automated error models and their potential to improve the efficiency of the data annotation process, specifically with respect to the quality and turnaround of annotations. We further analyze how automated error detection models can improve related processes such as auditing and relabeling. Compared to previous work addressing annotation errors, our approach introduces a unique emphasis on task-agnostic error prediction. Unlike many prior models, which focus on specific tasks and require separate training for each labeling task, our models are trained to predict errors directly from a combination of task features and behavioral features derived from the annotation process itself, offering a higher degree of generalizability.
@@ 1.1 Background and Related Work
Automated error detection techniques aim to identify instances that may be labeled incorrectly or inconsistently within a given dataset [klie_annotation_2022]. Once these potential errors are detected, they can be forwarded for manual review by human annotators or processed via other annotation error correction methods. Error detection has been implemented across diverse domains in natural language processing [barnes_sentiment_2019, alt_tacred_2020, klie_annotation_2022], and image classification [northcutt_pervasive_2021, bernhardt_active_2022]. However, the logic of error modeling can be applied to a variety of supervised ML tasks.
@@
Previous literature has distinguished between variation-based, rule-based, or model-based approaches to error detection [klie_annotation_2022]. Variation-based approaches generally rely on inputs labeled by multiple annotators. For example, in settings where only one label can be correct, disagreement across annotators can indicate annotation errors. Annotator disagreement can be captured by computing variation metrics like label entropy across annotators, such that high entropy would be indicative of annotation errors [hollenstein_inconsistency_2016]. While variation-based error detection methods are easy to implement, one major disadvantage is that they are limited to crowd-sourced scenarios, where more than one annotator label is produced for each instance. Requiring multiple labels can introduce significant costs to annotation processes.
@@
Rule-based techniques, on the other hand, are useful when there are clear annotation guidelines that can be translated into formal rules. In such situations, annotations that are impossible given the set of rules would be flagged as incorrect. For example, errors can be flagged if an annotator picks a label that is not applicable to the given task. In language annotation tasks, n-grams that are lexically or morphologically impossible would fall into this category [kveton_semi-automatic_2002]. However, while rule-based approaches can prove to be highly effective, they are typically tailored to a very specific task, imposing limits on generalizability. Additionally, they only apply to a specific subset of errors, namely those that are the result of objective rule violations.
@@
Model-based error detection approaches typically utilize probabilistic classifiers designed to detect annotation errors. Past work has frequently relied on the models predicting an annotation label (typically a multi-class classification problem) and subsequently comparing the predicted label to a label provided by a human annotator. If there is disagreement between the human annotation and the model-generated annotation, the annotation can be marked as inconsistent and additionally directed to a human auditor [larson_inconsistencies_2020, van_halteren_detection_2000]. Alternatively, model confidence scores or predicted probabilities can provide an indication of task difficulty [amiri_spotting_2018]. Implicitly, these approaches rely on the assumption that models perform worse on instances that are more difficult for human annotators as well. Consequently, instances that show low model performance - either by deviation from human annotations or by low model confidence scores - would be more likely to contain annotation errors.
@@
However, it is important to note that a model’s confidence score on the labeling task itself may not always be a reliable indicator of a human annotator’s potential error. For instance, low model confidence could be due to the model’s limitations in representing complex or context-specific tasks rather than indicating an annotation error. Conversely, there may be cases where a model is highly confident and yet wrong due to biases in training data or overly simplistic decision boundaries. Therefore, while using model confidence scores can certainly be a useful tool in identifying potential annotation errors, it should be treated as part of a broader toolkit and used in conjunction with other methods. For a more comprehensive approach to error detection, it might be beneficial also to consider other factors such as the context of the task, the difficulty level of the task, and the reliability and expertise of the human annotators themselves. Relatedly, model-based approaches can be less effective in identifying systematic errors, where both the model and the human annotators consistently misinterpret certain types of instances. Such errors could go undetected as model predictions and human annotations would agree, despite being incorrect. Finally, approaches predicting labels may suffer from low generalizability, as a separate model needs to be trained for each specific labeling task, which can be computationally expensive and may not be feasible for all tasks, especially those requiring complex models.
@@
As an alternative approach explored in this paper, a model can be used to directly predict whether a human label is correct or not (typically a binary classification problem) without explicitly predicting the underlying label itself. The advantage of this approach is that it is often easier to predict the presence of an error compared to predicting the correct label. Relatedly, directly predicting the correctness of a label can be computationally more efficient, as it does not necessarily need to consider all possible label classes, which can be vast in certain tasks. Overall this approach makes it easier to train, evaluate, and maintain such models even when training data are limited. Additionally, error detection models can be set up to be task agnostic: a model predicting errors instead of labels may generalize to a wider range of tasks because certain features are applicable to many different scenarios. For example, taking into account general measures of task complexity, past annotator performance, time on task, or comments provided by annotators can be informative irrespective of the underlying labeling task. Past research has rarely applied this approach [klie_annotation_2022] since most studies are based on archival data for which no annotator data (e.g., past performance) or task-completion data (e.g., time on task) are available.
@@ 1.2 Current Research
The current research presents a study on generalizable error detection models, taking into account behavioral features in addition to task features, and their applications for search relevance annotation tasks. The tasks were drawn from an extensive data annotation program that generates training and evaluation data for industry-scale search and recommendation systems. We focus on search relevance annotations related to music streaming, video streaming, and mobile applications. A detailed description of the annotation task can be found in the Method section below.
Drawing on authentic data from the search relevance annotation program, we first use model-based error detection techniques to show that annotation errors can be detected with moderate classification performance, based on a combination of task features and behavioral features derived from the data annotation process. We demonstrate the predictability of errors in search relevance rating tasks for all three ML applications: music streaming, video streaming, and mobile applications. Second, using Shapley values [lundberg_unified_2017], we explore the contributions of a variety of behavioral and task-related features in order to better understand which sources of information are driving the model’s predictions. Third, we explore the generalizability of our error detection models across rating tasks and ML applications. Fourth, we show that the error detection models can unlock considerable efficiency gains for auditing and relabeling by prioritizing tasks with high predicted error probabilities at the expense of tasks with low predicted error probabilities. Taken together, our findings show that behavioral error detection models can be highly effective in improving the efficiency of data annotation processes.
@@ 2.2 Search Relevance Annotation Tasks
{xtabular*}
p6.8cmp9cm Feature Name Description
@@
   input_query_type Intent classification for query
@@
   input_language Categorical variable distinguishing between 21 input languages
   storefront_name Categorical variable distinguishing between 23 geographic regions